OpenStudy (anonymous):

f(x) = 2x + 5 and g(x) = x − 7, solve for f(g(x)) when x = −3

OpenStudy (anonymous):

plug in the - 3 in g g(-3 ) = - 3 - 7 = - 10 then plug in -10 in f(x) f(-10) = 2 ( -10 ) + 5 = - 20 + 5 = 15

OpenStudy (anonymous):

i mean = - 15 not 15 x)

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

wlcm